OpenStudy (anonymous):

Find the sum or difference. 7/8+9/16 A 2/3 B 1 7/16 C 1 1/2

OpenStudy (anonymous):

b

OpenStudy (sleepyjess):

@amysparkly12 Please do not just give out answers.

OpenStudy (anonymous):

7/8+9/16= 23/16 which is not among your choices. Are you sure you typed the problem correct?

OpenStudy (sleepyjess):

Yes it is in the choices, it is just listed as \(1\dfrac{7}{16}\)

OpenStudy (anonymous):

Cool